# Learning Representations for Independence Testing

### v2

@@ 7 Discussion
We find that MMD-D underperforms HSIC-D in almost all settings, despite the equivalence between MMD and HSIC specified in Proposition 3.1. This equivalence, while true at the population level, is not the case when considering finite sample estimates. These estimators exhibit different statistical properties, with some being more conducive to learning than others. In particular, we notice that MMD estimators based on a single permutation of the samples (as described in Section 3.2) exhibit larger variance compared to HSIC ones. We prove this for the biased estimators in Section E.4, and observe the same phenomenon when training MMD-D and HSIC-D.
@@
We hypothesized that maximizing an estimate of the asymptotic power would produce stronger tests than maximizing the test statistic directly. Although this seems to be true for HSIC (Section E.3), it is surprisingly not the case for MI-based statistics; The previous power versus test size experiments show that directly maximizing InfoNCE or NWJ outperforms maximizing the NDS asymptotic power, at least when using a permutation test. Why is this?
Figure 6 examines several power estimates for an NDS test based on an InfoNCE-learned critic. Notice that the asymptotic formula (blue) agrees well with the simulated test power (orange), indicating that the asymptotics describe the test well. However, the permutation power (green) is significantly greater than what the asymptotics suggest. Recalling that the simulated test power is in fact using a near-exact threshold based on the empirical quantile of the test statistic under the null, this should be roughly the most powerful valid test with a sample-independent threshold. Permutation tests, however, use a sample-dependent threshold, which in this case apparently can yield a much more powerful test than the best sample-independent threshold. We explore this discrepancy further in Section E.7.
@@
[…]
@@ Appendix A Asymptotics of HSIC estimators
[…]
@@
Under the null hypothesis $\mathfrak{H}_{0}:\mathbb{P}_{xy}=\mathbb{P}_{x}\times\mathbb{P}_{y}$, the U-statistic estimator of $\operatorname{HSIC}$ is degenerate. In this case $m\widehat{\operatorname{HSIC}}_{\rm u}$ converges in distribution to
[…]
@@
where $h_{ijqr}:=h(i,j,q,r)$ is defined in Proposition A.1, and $F_{i,q,r}$ denotes the probability measure with respect to variables $z_{i}$, $z_{q}$, and $z_{r}$.
Appendix B Uniform Convergence: NDS
B.1 Preliminaries
@@
We note that $\hat{T}_{1}$ is just the NDS (2), and that $\hat{T}_{0}$ is the permutation invariant term in the numerator of Equation 12. For brevity, we often omit the subscripts $0,1$ and/or the parameters $\theta$ when it is uninformative or clear from the context.
@@ B.2 Assumptions
B.3 Lemmas
Lemma B.1 (Concentration of $\hat{T}_{1}$).
For any critic $f$ satisfying Assumption (C), with probability at least $1-\delta$ we have
@@ B.3 Lemmas
Lemma B.2 (Concentration of $\hat{T}_{0}$).
For any critic $f$ satisfying Assumption (C), with probability at least $1-\delta$ we have
@@
Lemma B.3 (Bias of $\hat{T}_{0}$).
For any critic $f$ satisfying Assumption (C), the bias of $\hat{T}_{0}$ is bounded by
@@
Lemma B.4 (Concentration of $\hat{\sigma}^{2}$).
For any critic $f$ satisfying Assumption (C), with probability at least $1-\delta$ we have
@@
where we use the finite differences computed in the proof of Lemma B.1 to bound $|\hat{T}_{1,(k)}-\hat{T}_{1}|$. Thus, with probability at least $1-\delta$,
@@
Lemma B.5 (Bias of $\hat{\sigma}^{2}$).
For any critic $f$ satisfying Assumption (C), the bias of $\hat{\sigma}^{2}$ is bounded by
@@
Lemma B.6 (Concentration of $\hat{\xi}^{2}$).
For any critic $f$ satisfying Assumption (C), with probability at least $1-\delta$ we have
@@
$\sup_{\begin{subarray}{c}x^{\prime},y^{\prime}\\ \mathbb{X},\mathbb{Y}\end{subarray}}|\hat{\xi}_{(\ell)}^{2}-\hat{\xi}|$ $\leq\sup_{\begin{subarray}{c}x^{\prime},y^{\prime}\\ \mathbb{X},\mathbb{Y}\end{subarray}}|\hat{\nu}_{2,(\ell)}-\hat{\nu}_{2}|+\sup_{\begin{subarray}{c}x^{\prime},y^{\prime}\\ \mathbb{X},\mathbb{Y}\end{subarray}}|\hat{T}_{0,(\ell)}^{2}-\hat{T}_{0}^{2}|$
@@
Lemma B.7 (Bias of $\hat{\xi}^{2}$).
For any critic $f$ satisfying Assumption (C), the bias of $\hat{\xi}^{2}$ is bounded by
@@
B.4 Propositions
Proposition B.8 (Uniform convergence of $\hat{T}_{1}$).
For any critic family $\{f_{\theta}:\theta\in\Theta\}$ and parameter space $\Theta$ satisfying Assumptions (A), (B) and (C), we have with probability at least $1-\delta$ that
@@ B.4 Propositions
We employ a covering-number argument on the parameter space. Let $\mathcal{U}=\{\theta_{i}\}_{i=1..N}\subseteq\Theta$ be a $\rho$-cover of $\Theta$; i.e., any point in $\Theta$ is in the closed $\rho$-neighborhood of some point in $\mathcal{U}$. Assumption (A) ensures the minimal cover exists with at most $N=(4R/\rho)^{D}$ points (Cucker & Smale, 2002, Proposition 5). Now, for any $\theta\in\Theta$, let $\theta^{\prime}$ be the center in $\mathcal{U}$ closest to $\theta$, and then decompose the bound into
@@
We first handle terms (a) and (c). From Assumption (B), we have
@@
Next, we find a high probability bound on (b). Recalling the result of Lemma B.1, and then taking the union bound over all $N$ points in $\mathcal{U}$, we get that with probability at least $1-\delta$
@@
Proposition B.9 (Uniform convergence of $\hat{T}_{0}$).
For any critic family $\{f_{\theta}:\theta\in\Theta\}$ and parameter space $\Theta$ satisfying Assumptions (A), (B) and (C), we have with probability at least $1-\delta$ that
@@
We employ a covering-number argument on the parameter space. Let $\mathcal{U}=\{\theta_{i}\}_{i=1..N}\subseteq\Theta$ be a $\rho$-cover of $\Theta$; i.e., any point in $\Theta$ is in the closed $\rho$-neighborhood of some point in $\mathcal{U}$. Assumption (A) ensures the minimal cover exists with at most $N=(4R/\rho)^{D}$ points (Cucker & Smale, 2002, Proposition 5). Now, for any $\theta\in\Theta$, let $\theta^{\prime}$ be the center in $\mathcal{U}$ closest to $\theta$, and then decompose the bound into
@@
We first handle terms (a) and (d). From Assumption (B), we have
@@
Next, we find a high probability bound on (b). Recalling the result of Lemma B.2, and then taking the union bound over all $N$ points in $\mathcal{U}$, we get that with probability at least $1-\delta$
@@
Finally, combining everything and using Lemma B.3 to bound (c), we get that with probability at least $1-\delta$
@@
Proposition B.10 (Uniform convergence of $\hat{\sigma}^{2}$).
For any critic family $\{f_{\theta}:\theta\in\Theta\}$ and parameter space $\Theta$ satisfying Assumptions (A), (B) and (C), we have with probability at least $1-\delta$ that
@@
We employ a covering-number argument on the parameter space. Let $\mathcal{U}=\{\theta_{i}\}_{i=1..N}\subseteq\Theta$ be a $\rho$-cover of $\Theta$; i.e., any point in $\Theta$ is in the closed $\rho$-neighborhood of some point in $\mathcal{U}$. Assumption (A) ensures the minimal cover exists with at most $N=(4R/\rho)^{D}$ points (Cucker & Smale, 2002, Proposition 5). Now, for any $\theta\in\Theta$, let $\theta^{\prime}$ be the center in $\mathcal{U}$ closest to $\theta$, and then decompose the bound into
@@
We first handle terms (a) and (d). From Assumption (B), we have
@@
$\quad+\sup_{\theta\in\Theta}\operatorname*{\mathbb{E}}\Big{|}f_{\theta^{\prime}}(X,Y)-f_{\theta}(X,Y)\Big{|}\operatorname*{\mathbb{E}}\Big{|}f_{\theta^{\prime}}(\tilde{X},\tilde{Y})+f_{\theta}(\tilde{X},\tilde{Y})\Big{|}$
@@
Next, we find a high probability bound on (b). Recalling the result of Lemma B.4, and then taking the union bound over all $N$ points in $\mathcal{U}$, we get that with probability at least $1-\delta$
@@
Finally, combining (a), (b), (d), and using Lemma B.5 to bound (c), we get that with probability at least $1-\delta$
@@
Proposition B.11 (Uniform convergence of $\hat{\xi}^{2}$).
For any critic family $\{f_{\theta}:\theta\in\Theta\}$ and parameter space $\Theta$ satisfying Assumptions (A), (B) and (C), we have with probability at least $1-\delta$ that
@@
We employ a covering-number argument on the parameter space. Let $\mathcal{U}=\{\theta_{i}\}_{i=1..N}\subseteq\Theta$ be a $\rho$-cover of $\Theta$; i.e., any point in $\Theta$ is in the closed $\rho$-neighborhood of some point in $\mathcal{U}$. Assumption (A) ensures the minimal cover exists with at most $N=(4R/\rho)^{D}$ points (Cucker & Smale, 2002, Proposition 5). Now, for any $\theta\in\Theta$, let $\theta^{\prime}$ be the center in $\mathcal{U}$ closest to $\theta$, and then decompose the bound into
@@
First, we bound terms (a) and (d) using Assumption (B) and the covering-number argument. Let $T_{\theta}=\operatorname*{\mathbb{E}}[f_{\theta}(X,Y)]$ and $\nu_{\theta}=\operatorname*{\mathbb{E}}[f_{\theta}(X,Y)^{2}]$ be the first and second moments under the null distribution, so that $\xi_{\theta}^{2}=\nu_{\theta}-T_{\theta}^{2}$. It follows that
@@
Next, we bound (b) using Lemma B.6 and a uniform bound over all $N$ points in the $\rho$-cover. We get that with probability at least $1-\delta$
@@
Finally, combining (a), (b), (d), and using Lemma B.7 to bound (c), we get that with probability at least $1-\delta$
@@
B.5 Main Results
Theorem B.12 (Uniform convergence of $\hat{J}_{\text{NDS}}^{\lambda}$).
Let $\{f_{\theta}:\theta\in\Theta\}$ be a critic family with parameter space $\Theta$ satisfying Assumptions (A), (B) and (C). Suppose $\sigma_{\theta}^{2}\geq s^{2}$ and $\xi_{\theta}^{2}\geq s^{2}$ for some positive $s$, and $\lambda=n^{-1/3}$. Then, with probability at least $1-\delta$,

### v3

@@ 7 Discussion
We find that MMD-D underperforms HSIC-D in almost all settings, despite the equivalence between MMD and HSIC specified in Proposition˜3.1. This equivalence, while true at the population level, is not the case when considering finite sample estimates. These estimators exhibit different statistical properties, with some being more conducive to learning than others. In particular, we notice that MMD estimators based on a single permutation of the samples (as described in Section˜3.2) exhibit larger variance compared to HSIC ones. We prove this for the biased estimators in Section˜F.4, and observe the same phenomenon when training MMD-D and HSIC-D.
@@
We hypothesized that maximizing an estimate of the asymptotic power would produce stronger tests than maximizing the test statistic directly. Although this seems to be true for HSIC (Section˜F.3), it is surprisingly not the case for MI-based statistics; The previous power versus test size experiments show that directly maximizing InfoNCE or NWJ outperforms maximizing the NDS asymptotic power, at least when using a permutation test. Why is this?
Figure˜6 examines several power estimates for an NDS test based on an InfoNCE-learned critic. Notice that the asymptotic formula (blue) agrees well with the simulated test power (orange), indicating that the asymptotics describe the test well. However, the permutation power (green) is significantly greater than what the asymptotics suggest. Recalling that the simulated test power is in fact using a near-exact threshold based on the empirical quantile of the test statistic under the null, this should be roughly the most powerful valid test with a sample-independent threshold. Permutation tests, however, use a sample-dependent threshold, which in this case apparently can yield a much more powerful test than the best sample-independent threshold. We explore this discrepancy further in Section˜F.7.
@@
Figure 6: Power of NDS tests derived from a learned InfoNCE critic $f$. For each problem, the blue line (asymp) is power described by the asymptotic formula (3), where we use a very large sample size (n=20 000n=20\,000) to estimate the population level statistics $T,T_{0}$ and their deviations $\sigma_{\mathfrak{H}_{1}},\sigma_{\mathfrak{H}_{0}}$. The orange line (empirical) computes the power based on the simulated null distribution, i.e. the threshold estimated via the empirical CDF of $\hat{T}_{0}$ with test size $m$. The green line (perm) is the permutation test power.
@@ Appendix A Asymptotics of HSIC Estimators
$\sqrt{m}(\widehat{\operatorname{HSIC}}_{\rm u}-\operatorname{HSIC})\overset{\textbf{d}}{\to}\mathcal{N}(0,\sigma^{2}_{\mathfrak{H}_{1}}).$ (14)
@@
Under the null hypothesis $\mathfrak{H}_{0}:\mathbb{P}_{xy}=\mathbb{P}_{x}\times\mathbb{P}_{y}$, the U-statistic estimator of $\operatorname{HSIC}$ is degenerate. In this case $m\widehat{\operatorname{HSIC}}_{\rm u}$ converges in distribution as
$m\widehat{\operatorname{HSIC}}_{\rm u}\overset{\textbf{d}}{\to}\sum_{\ell=1}^{\infty}\lambda_{\ell}(z^{2}_{\ell}-1),$
@@
where $h_{ijqr}:=h(i,j,q,r)$ is defined in Proposition˜A.1, and $F_{i,q,r}$ denotes the probability measure with respect to variables $z_{i}$, $z_{q}$, and $z_{r}$.
Appendix B Bernstein-based Justification for the SNR
For both the NDS and HSIC tests, our justifications of the SNR criterion for test power, (3) and (6), relied on the asymptotic central limit theorem. A Berry-Esseen theorem would provide uniform bounds on the $o(1)$ terms that appear in these derivations, but one might expect for HSIC in particular that these might not be especially tight in some cases: since the null distribution is asymptotically a mixture of chi-squared variables (Proposition˜A.2), when the dependence under the current kernel is weak, the distribution may be further from normal.
We can also justify the same criterion, although less precisely, with a non-asymptotic argument. While we present it here for HSIC, it also applies in the same way to NDS. The Bernstein-type bound of Maurer (2017, Theorem 6) gives that, for kernels bounded as in Assumption˜(C’),
$\Pr\left(\lvert\widehat{\operatorname{HSIC}}-\operatorname{HSIC}\rvert>\varepsilon\right)\leq 2\exp\left(\frac{-n\varepsilon^{2}}{2\sigma_{\mathfrak{H}_{1}}^{2}+\frac{2304\nu_{k}\nu_{l}}{n-4}+\frac{256}{3}\varepsilon}\right)\leq\delta.$
Solving for $\varepsilon$, we obtain
$\frac{-m\varepsilon^{2}}{2\sigma_{\mathfrak{H}_{1}}^{2}+\frac{2304\nu_{k}\nu_{l}}{m-4}+\frac{256}{3}\varepsilon}\leq\log\frac{\delta}{2}$
$\varepsilon^{2}\geq\frac{1}{m}\log\frac{2}{\delta}\left(2\sigma_{\mathfrak{H}_{1}}^{2}+\frac{2304\nu_{k}\nu_{l}}{m-4}+\frac{256}{3}\varepsilon\right)$
$\varepsilon^{2}-\frac{256}{3m}\log\frac{2}{\delta}\varepsilon-\left(\frac{2}{m}\sigma_{\mathfrak{H}_{1}}^{2}\log\frac{2}{\delta}+\frac{2304\nu_{k}\nu_{l}}{m-4}\right)\geq 0.$
This can be written $\varepsilon^{2}-B\varepsilon-C^{2}\geq 0$ for positive $B,C$; restricting to positive $\varepsilon$, this occurs when $\varepsilon\geq\frac{1}{2}B+\frac{1}{2}\sqrt{B^{2}+4C^{2}}$, and hence also occurs when $\varepsilon$ is larger than the (slightly larger but simpler) bound $B+C$. That is, with probability at least $1-\delta$, it holds that
$\lvert\widehat{\operatorname{HSIC}}-\operatorname{HSIC}\rvert$ $\leq\frac{256}{3m}\log\frac{2}{\delta}+\sqrt{\frac{2}{m}\sigma_{\mathfrak{H}_{1}}^{2}\log\frac{2}{\delta}+\frac{2304\nu_{k}\nu_{l}}{m-4}}$
$\leq\frac{256}{3m}\log\frac{2}{\delta}+\sigma_{\mathfrak{H}_{1}}\sqrt{\frac{2}{m}\log\frac{2}{\delta}}+\sqrt{\frac{2304\nu_{k}\nu_{l}}{m-4}}.$
Recall from the discussion preceding (6) that the rejection threshold $r_{m}$ is $\frac{1}{m}\left(\Psi^{-1}(1-\alpha)+o(1)\right)$ (where $\Psi$ will vary with the kernel choice). Now, a sufficient condition for the test rejecting is that both $\widehat{\operatorname{HSIC}}\geq\operatorname{HSIC}-\varepsilon$ and $\operatorname{HSIC}-\varepsilon>r_{m}$. Plugging in the deviation from above, we obtain power at least $1-\delta$ as long as
$\operatorname{HSIC}$ $>r_{m}+\frac{256}{3m}\log\frac{2}{\delta}+\sigma_{\mathfrak{H}_{1}}\sqrt{\frac{2}{m}\log\frac{2}{\delta}}+\sqrt{\frac{2304\nu_{k}\nu_{l}}{m-4}}$
$=\sigma_{\mathfrak{H}_{1}}\sqrt{\frac{2}{m}\log\frac{2}{\delta}}+\frac{1}{m}\left(\Psi^{-1}(1-\alpha)+\operatorname{\mathcal{O}}(1)\right)+48\sqrt{\frac{\nu_{k}\nu_{l}}{m-4}},$
absorbing $\frac{256}{3m}\log\frac{2}{\delta}=\operatorname{\mathcal{O}}(1/m)$. Equivalently, the power condition holds if
$\frac{\operatorname{HSIC}}{\sigma_{\mathfrak{H}_{1}}}-\frac{48}{\sqrt{m-4}}\frac{\sqrt{\nu_{k}\nu_{l}}}{\sigma_{\mathfrak{H}_{1}}}>\sqrt{\frac{2}{m}\log\frac{2}{\delta}}+\frac{1}{m}\frac{\Psi^{-1}(1-\alpha)}{\sigma_{\mathfrak{H}_{1}}}+\operatorname{\mathcal{O}}\left(\frac{1}{m\sigma_{\mathfrak{H}_{1}}}\right).$
Now, we know that scaling $k$ or $l$ should not change the difficulty of the problem: indeed, it will scale $\operatorname{HSIC}$, $\sigma_{\mathfrak{H}_{1}}$, $\Psi^{-1}(1-\alpha)$, and $\sqrt{\nu_{k}\nu_{l}}$ all equally. Thus, the only term in this inequality which is potentially sensitive to scaling is the final $\operatorname{\mathcal{O}}$ term. From this point, then, while other changes in the kernel could potentially affect $\Psi^{-1}(1-\alpha)/\sigma_{\mathfrak{H}_{1}}$ in meaningful ways, for reasonably large $m$ and $\sigma_{\mathfrak{H}_{1}}$ not too drastically small, the right-hand side is only weakly dependent on the kernel choice. It is thus reasonable to think that we can maximize the left-hand side of the inequality, and ignore the right-hand side, to roughly maximize power. It seems highly unlikely, however, for the coefficient $48/\sqrt{m-4}$ to be exactly the right trade-off between $\operatorname{HSIC}$ and the kernel bound; this comes out of the details of the (conservative and somewhat loose) concentration inequalities to reach this point. To maximize our chance at having high power, then, it is very reasonable to maximize the SNR.
Appendix C Uniform Convergence: NDS
C.1 Preliminaries
@@
We note that $\hat{T}_{1}$ is just the NDS (2), and that $\hat{T}_{0}$ is the permutation invariant term in the numerator of Equation˜12. For brevity, we often omit the subscripts $0,1$ and/or the parameters $\theta$ when it is uninformative or clear from the context.
@@ C.2 Assumptions
C.3 Lemmas
Lemma C.1 (Concentration of $\hat{T}_{1}$).
For any critic $f$ satisfying Assumption˜(C), with probability at least $1-\delta$ we have
@@ C.3 Lemmas
Lemma C.2 (Concentration of $\hat{T}_{0}$).
For any critic $f$ satisfying Assumption˜(C), with probability at least $1-\delta$ we have
@@
Lemma C.3 (Bias of $\hat{T}_{0}$).
For any critic $f$ satisfying Assumption˜(C), the bias of $\hat{T}_{0}$ is bounded by
@@
Lemma C.4 (Concentration of $\hat{\sigma}^{2}$).
For any critic $f$ satisfying Assumption˜(C), with probability at least $1-\delta$ we have
@@
where we use the finite differences computed in the proof of Lemma˜C.1 to bound $|\hat{T}_{1,(k)}-\hat{T}_{1}|$. Thus, with probability at least $1-\delta$,
@@
Lemma C.5 (Bias of $\hat{\sigma}^{2}$).
For any critic $f$ satisfying Assumption˜(C), the bias of $\hat{\sigma}^{2}$ is bounded by
@@
Lemma C.6 (Concentration of $\hat{\xi}^{2}$).
For any critic $f$ satisfying Assumption˜(C), with probability at least $1-\delta$ we have
@@
[…]
@@
Lemma C.7 (Bias of $\hat{\xi}^{2}$).
For any critic $f$ satisfying Assumption˜(C), the bias of $\hat{\xi}^{2}$ is bounded by
@@
C.4 Propositions
Proposition C.8 (Uniform convergence of $\hat{T}_{1}$).
For any critic family $\{f_{\theta}:\theta\in\Theta\}$ and parameter space $\Theta$ satisfying Assumptions˜(A), (B) and (C), we have with probability at least $1-\delta$ that
@@ C.4 Propositions
We employ a covering-number argument on the parameter space. Let $\mathcal{U}=\{\theta_{i}\}_{i=1..N}\subseteq\Theta$ be a $\rho$-cover of $\Theta$; i.e., any point in $\Theta$ is in the closed $\rho$-neighborhood of some point in $\mathcal{U}$. Assumption˜(A) ensures the minimal cover exists with at most $N=(4R/\rho)^{D}$ points (Cucker & Smale, 2002, Proposition 5). Now, for any $\theta\in\Theta$, let $\theta^{\prime}$ be the center in $\mathcal{U}$ closest to $\theta$, and then decompose the bound into
@@
We first handle terms (a) and (c). From Assumption˜(B), we have
@@
Next, we find a high probability bound on (b). Recalling the result of Lemma˜C.1, and then taking the union bound over all $N$ points in $\mathcal{U}$, we get that with probability at least $1-\delta$
@@
Proposition C.9 (Uniform convergence of $\hat{T}_{0}$).
For any critic family $\{f_{\theta}:\theta\in\Theta\}$ and parameter space $\Theta$ satisfying Assumptions˜(A), (B) and (C), we have with probability at least $1-\delta$ that
@@
We employ a covering-number argument on the parameter space. Let $\mathcal{U}=\{\theta_{i}\}_{i=1..N}\subseteq\Theta$ be a $\rho$-cover of $\Theta$; i.e., any point in $\Theta$ is in the closed $\rho$-neighborhood of some point in $\mathcal{U}$. Assumption˜(A) ensures the minimal cover exists with at most $N=(4R/\rho)^{D}$ points (Cucker & Smale, 2002, Proposition 5). Now, for any $\theta\in\Theta$, let $\theta^{\prime}$ be the center in $\mathcal{U}$ closest to $\theta$, and then decompose the bound into
@@
We first handle terms (a) and (d). From Assumption˜(B), we have
@@
Next, we find a high probability bound on (b). Recalling the result of Lemma˜C.2, and then taking the union bound over all $N$ points in $\mathcal{U}$, we get that with probability at least $1-\delta$
@@
Finally, combining everything and using Lemma˜C.3 to bound (c), we get that with probability at least $1-\delta$
@@
Proposition C.10 (Uniform convergence of $\hat{\sigma}^{2}$).
For any critic family $\{f_{\theta}:\theta\in\Theta\}$ and parameter space $\Theta$ satisfying Assumptions˜(A), (B) and (C), we have with probability at least $1-\delta$ that
@@
We employ a covering-number argument on the parameter space. Let $\mathcal{U}=\{\theta_{i}\}_{i=1..N}\subseteq\Theta$ be a $\rho$-cover of $\Theta$; i.e., any point in $\Theta$ is in the closed $\rho$-neighborhood of some point in $\mathcal{U}$. Assumption˜(A) ensures the minimal cover exists with at most $N=(4R/\rho)^{D}$ points (Cucker & Smale, 2002, Proposition 5). Now, for any $\theta\in\Theta$, let $\theta^{\prime}$ be the center in $\mathcal{U}$ closest to $\theta$, and then decompose the bound into
@@
We first handle terms (a) and (d). From Assumption˜(B), we have
@@
[…]
@@
Next, we find a high probability bound on (b). Recalling the result of Lemma˜C.4, and then taking the union bound over all $N$ points in $\mathcal{U}$, we get that with probability at least $1-\delta$
@@
Finally, combining (a), (b), (d), and using Lemma˜C.5 to bound (c), we get that with probability at least $1-\delta$
@@
Proposition C.11 (Uniform convergence of $\hat{\xi}^{2}$).
For any critic family $\{f_{\theta}:\theta\in\Theta\}$ and parameter space $\Theta$ satisfying Assumptions˜(A), (B) and (C), we have with probability at least $1-\delta$ that
@@
We employ a covering-number argument on the parameter space. Let $\mathcal{U}=\{\theta_{i}\}_{i=1..N}\subseteq\Theta$ be a $\rho$-cover of $\Theta$; i.e., any point in $\Theta$ is in the closed $\rho$-neighborhood of some point in $\mathcal{U}$. Assumption˜(A) ensures the minimal cover exists with at most $N=(4R/\rho)^{D}$ points (Cucker & Smale, 2002, Proposition 5). Now, for any $\theta\in\Theta$, let $\theta^{\prime}$ be the center in $\mathcal{U}$ closest to $\theta$, and then decompose the bound into
@@
First, we bound terms (a) and (d) using Assumption˜(B) and the covering-number argument. Let $T_{\theta}=\operatorname*{\mathbb{E}}[f_{\theta}(X,Y)]$ and $\nu_{\theta}=\operatorname*{\mathbb{E}}[f_{\theta}(X,Y)^{2}]$ be the first and second moments under the null distribution, so that $\xi_{\theta}^{2}=\nu_{\theta}-T_{\theta}^{2}$. It follows that
@@
Next, we bound (b) using Lemma˜C.6 and a uniform bound over all $N$ points in the $\rho$-cover. We get that with probability at least $1-\delta$
@@
Finally, combining (a), (b), (d), and using Lemma˜C.7 to bound (c), we get that with probability at least $1-\delta$
@@
C.5 Main Results
Theorem C.12 (Uniform convergence of $\hat{J}_{\text{NDS}}^{\lambda}$).
Let $\{f_{\theta}:\theta\in\Theta\}$ be a critic family with parameter space $\Theta$ satisfying Assumptions˜(A), (B) and (C). Suppose $\sigma_{\theta}^{2}\geq s^{2}$ and $\xi_{\theta}^{2}\geq s^{2}$ for some positive $s$, and $\lambda=n^{-1/3}$. Then, with probability at least $1-\delta$,